OpenStudy (diana.xl):

A new online store charges all customers one price to download any song and another price to download any game. Jenna paid $20.50 to download eight songs and three games, and Dwayne paid $22.50 to download four songs and five games. How much does the store charge to download each song and game?

OpenStudy (diana.xl):

a. Write a system of linear equations based on the description. Use x to represent the cost of each song and y to represent the cost of each game. b. Solve the system using the elimination method. Show all the steps leading to your answer.

OpenStudy (diana.xl):

@imqwerty

OpenStudy (diana.xl):

@ganeshie8

OpenStudy (marigirl):

Let's denote Songs: s Games: g Jenna buys 8 songs and 3 games 8s+3g=20.50 Dwayne 4s+5g=22.50 With me so far

OpenStudy (marigirl):

Opps didn't read ur comment 8x+3y=20.50 4x+5y=22.50

OpenStudy (marigirl):

8x+3y=20.45 4x+5y=22.50 ( multiply by two) This will allow us to eliminate a variable so we can find the value of other variable 8x+3y=20.50 8x+10y=45 - _____________ -7y =--24.50 y=3.5 So a game costs $3.50

OpenStudy (marigirl):

To find the value of a song, use any if the two equations. 8x+3y=20.50 We know now y=3.50 3 games will cost 10.50 8x+10.5=20.5 8x=10 X=1.25 A song costs $1.25